OpenStudy (anonymous):

what is the answer to x cubed +4x squared -3x-12 over x squared -3 ?

OpenStudy (mertsj):

\[\frac{x^3+4x^2-3x-12}{x^2-3}=\frac{(x^2-3)(x+4)}{x^2-3}\]

OpenStudy (mertsj):

Cancel the common factor.